OpenStudy (anonymous):

solve equation 1/3x=1/10x+10/3

OpenStudy (anonymous):

multiply all 3 terms by the lowest common denominator (30x) 30x(1/3x) = 30x(1/10x) + 30x(10/3) 10 = 3 +100x 7=100x x=.07

OpenStudy (anonymous):

answers are -10/3,7/100,0,0r 100/7

OpenStudy (anonymous):

.07 = 7/100

OpenStudy (anonymous):

Is the equation 1 over 3x or one third x? 1 over 10x or one tenth x?